A union charges monthly dues of $6.00 plus $0.15 for each hour worked during the month. A union member's dues for March were $31.80. Write an equation to find the number of hours, x, the union member worked during the month of March.
How many hours did the union member work during the month of March?
x = h?

Answer by user_dude2008(1862)

0.15x+6=31.80

0.15x=31.80-6
0.15x=25.80

x=25.80/0.15

x=172

Answer: 172 hours